A car and a bus leave town at 1 P.M. and head for a town 300 mi away. The rate of the car is twice the rate of the bus. The car arrives 5 h ahead of the bus. Find the rate of the car.
Answer by stanbon(75887)

A car and a bus leave town at 1 P.M. and head for a town 300 mi away. The rate of the car is twice the rate of the bus. The car arrives 5 h ahead of the bus. Find the rate of the car.
---
Car DATA:
distance = 300 mi ; rate = 2x mph ; time = d/r = 300/(2x) hrs
-------------
Bus DATA:
distance = 300 mi ; rate = x mph ; time = d/r = 300/x hrs
--------------------------------------------------
Equation:
Bus time - Car time = 5 hrs.
300/x - 300/(2x) = 5
600 - 300 = 10x
10x = 300
x = 30 mph
2x = 60 mph (rate of the car)
==================================
Cheers,
Stan H.